OpenStudy (anonymous):

What is |-8| + |8-2| ?

OpenStudy (chaise):

14. Let me know if you want to know how to solve questions like these.

OpenStudy (anonymous):

8+8-2=8+6=14

OpenStudy (anonymous):

modulus is always positive

OpenStudy (saifoo.khan):

8 + 6